OpenStudy (anonymous):

Please please help!! The first four terms of a sequence are shown below: 6, 2, -2, -6 Which of the following functions best defines this sequence? f(1) = 6, f(n + 1) = f(n) + 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 2; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 6, f(n + 1) = f(n) + 2; for n ≥ 1

jimthompson5910 (jim_thompson5910):

Ignore the answer choices what's the rule for "6, 2, -2, -6"

OpenStudy (anonymous):

-4?

jimthompson5910 (jim_thompson5910):

be more specific

OpenStudy (anonymous):

How so?

jimthompson5910 (jim_thompson5910):

what do you mean by "-4"

OpenStudy (anonymous):

Oh well if you subtract four from six you get two. Four from two you get negative two and so on

jimthompson5910 (jim_thompson5910):

good, or if you add -4 to each term, you get the next term

jimthompson5910 (jim_thompson5910):

f(n+1) is the next term after f(n) which is why the rule is f(n+1) = f(n) + (-4) or f(n+1) = f(n) - 4

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np